OpenStudy (anonymous):

I'm completely lost with this problem, I just keep going in circles. I need to solve the problem by completing the square and factoring. Any help please? Y=-.6x^2+4.5x-2.6875

OpenStudy (anonymous):

do you mean "solve" as in set equal to zero and solve? or do you mean "complete the square" by writing as \(y=a(x-h)^2+k\)?

OpenStudy (anonymous):

guess i'll never know

OpenStudy (anonymous):

Solve as in find the roots using both of these methods, I all ready know the roots are .654 and 6.84

OpenStudy (anonymous):

so you want so solve \[-.6x^2+4.5x-2.6875=0\] by completing the square not so bad, especially with a calculator first step is to divide everything by \(-.6\) because in order to complete the square, the leading coefficient has to be 1

OpenStudy (anonymous):

\[x^2-\frac{15}{2}x+\frac{215}{48}=0\] then \[x^2-\frac{15}{2}=-\frac{215}{48}\] \[(x-\frac{15}{4})^2=\frac{215}{48}+(\frac{15}{4})^2\] and i will let you finish the arithmetic form here

OpenStudy (anonymous):

Thanks a bunch I got it :) Here's my work just in case you wanted to see what your tutoring did 0=x^2-15/2x+215/48 x^2-15/2X=-215/48 (X-15/4)^2=-214/48+225/16 (X-15/4)^2=115/12 X-15/4=+-SQRT(115/12) x=15/4+-SQRT(115/12)

OpenStudy (anonymous):

Now how do you do it by factoring?

OpenStudy (anonymous):

ok everything looks good, but i am not sure about the arithmetic

OpenStudy (anonymous):

I checked my answer, the roots checked out

OpenStudy (anonymous):

oh never mind, you had a typo and wrote 214, but it is 215 and you are correct ignore the last comment

OpenStudy (anonymous):

Oh ok sorry about that

OpenStudy (anonymous):

you can't really do it by factoring unless you know the roots since you know the roots you can claim it factors as \((x-r_1)(x-r_2)\)

OpenStudy (anonymous):

but really you could never come up with these numbers like you could for ones that factor using integers you have to know the answer to know how to factor

OpenStudy (anonymous):

Oh yes I see. Thanks a bunch for your help

OpenStudy (anonymous):

yw